In the lab, Justin has two solutions that contain alcohol and is mixing them with each other. He uses 3 times as much Solution A as Solution B. Solution A is 16% alcohol and Solution B is 11% alcohol. How many milliliters of Solution B does he use, if the resulting mixture has 177 milliliters of pure alcohol?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39617)

SOLUTION       PERCENT ALC.      QTY.        PURE ALC.

A                  16            3v

B                  11             v

                                              177

SOLUTION       PERCENT ALC.      QTY.        PURE ALC.

A                  16            3v         0.16*3v

B                  11             v         0.11v

                                              177

300

Answer by ikleyn(52781)

.
In the lab, Justin has two solutions that contain alcohol and is mixing them with each other.
He uses 3 times as much Solution A as Solution B. Solution A is 16% alcohol and Solution B
is 11% alcohol. How many milliliters of Solution B does he use, if the resulting mixture has
177 milliliters of pure alcohol?
~~~~~~~~~~~~~~~~~~~

Let x be the volume of solution B in milliliters (the quantity under the problem's question).

Then the volume of the solution A is 3x milliliters, according to the problem.


Solution B contains 0.11x milliliters of the pure alcohol.

Solution A contains 0.16*(3x) milliliters of the pure alcohol.


Write equation for the total content of the pure alcohol

    0.11x + 0.16*(3x) = 177  milliliter of the pure alcohol.


Simplify this equation and find x

    0.11x + 0.48x = 177

        0.59x     = 177

            x     = 177/0.59 = 300.


ANSWER.  Justin used 300 mL of solution B.

Solved.